OpenStudy (anonymous):

Alg 1 Honors! 1. What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−4, 0) and (−5, 0) (−4, 0) and (5, 0) (4, 0) and (−5, 0) (4, 0) and (5, 0)

OpenStudy (anonymous):

Oh I got the 1st one already:) I just need help with the second one

OpenStudy (ivettef365):

so you have 3x^2 +14x +8 (3x+2)(x+4) = 0

OpenStudy (anonymous):

Yupp. I need help with the 2nd one though

OpenStudy (ivettef365):

when y = 0 then you have 0 + 20 = x^2 - x 0 = x^2-x-20 now you can solve

OpenStudy (ivettef365):

you understand ?

OpenStudy (anonymous):

(x-5)(x+4) so (4, 0) and (−5, 0)

OpenStudy (ivettef365):

x-5 = 0 and x+4 = 0 x = 5 and x = -4 (5,0) and (-4,0)

OpenStudy (anonymous):

Ohhhh yeah I forgot to do that :) Thank you! Can you help me with another problem?

OpenStudy (anonymous):

What are the exact solutions of x2 − x − 4 = 0? a. x = the quantity of negative 1 plus or minus the square root of 15 all over 2 b. x = the quantity of 1 plus or minus the square root of 15 all over 2 c. x = the quantity of 1 plus or minus the square root of 17 all over 2 d. x = the quantity of negative 1 plus or minus the square root of 17 all over 2

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

use quadratic equation

OpenStudy (ivettef365):

did you get it??

OpenStudy (ivettef365):

??

OpenStudy (anonymous):

No i didnt :( Can you tell me the answer please?

OpenStudy (ivettef365):

did you try the website

OpenStudy (anonymous):

what do i put in the sqares?

OpenStudy (anonymous):

1, 1, and -4?

OpenStudy (anonymous):

@ivettef365 I got it Thank you! :)

OpenStudy (ivettef365):

I got c, hope you had the same :)